xx CarpetRegrid2

CarpetRegrid2

Erik Schnetter <schnetter@cct.lsu.edu>

February 15, 2006

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11activex1 10Scope: privatex01 BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex10  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex2  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex3  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex4  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex5  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex6  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex7  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex8  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11activex9  10Scope: privatex01  BOOLEAN

11Description: Is this region active?

11                       10  x01  Default: yes

11

11adaptivexblockxsize  10Scope: privatex01  INT

11Description: Block size for adaptive refinement

11  Range  10  x01  Default: 8

111:*  10

11adaptivexblockxsizexx  10Scope: privatex01  INT

11Description: Block size in x direction for adaptive refinement

11 Range 10 x01 Default: -1
11-1  10use adaptivexblockxsize

111:*  10

11

11adaptivexblockxsizexy  10Scope: privatex01  INT

11Description: Block size in y direction for adaptive refinement

11 Range 10 x01 Default: -1
11-1  10use adaptivexblockxsize

111:*  10

11

11adaptivexblockxsizexz  10Scope: privatex01  INT

11Description: Block size in z direction for adaptive refinement

11 Range 10 x01 Default: -1
11-1  10use adaptivexblockxsize

111:*  10

11

11adaptivexrefinement  10Scope: privatex01  BOOLEAN

11Description: Use levelxmask for adaptive refinement

11                       10  x01  Default: no

11

11addxlevelsxautomatically 10Scope: privatex01  BOOLEAN

11Description: Automatically add a new refinement level at each regrid

11                       10  x01  Default: no

11

11boundaryxshiftout  10Scope: privatex01  INT

11Description: Number of grid points added to the refinement boundary radius

11  Range  10  x01  Default: (none)

11*:*  10

11

11ensurexproperxnesting  10Scope: privatex01  BOOLEAN

11Description: Ensure proper nesting automatically

11                       10  x01  Default: yes

11

11expectxsymmetricxgrids 10Scope: privatex01  BOOLEAN

11Description: Expect a grid structure that is symmetric about the origin, and abort if it is not

11                       10  x01  Default: no

11

11freezexunalignedxlevels 10Scope: privatex01  BOOLEAN

11Description: Do not change refinement levels that do not exist at this time

11                       10  x01  Default: no

11

11freezexunalignedxparentxlevels 10Scope: privatex01  BOOLEAN

11Description: Do not change refinement levels where the parent does not exist at this time

11                       10  x01  Default: no

11

11granularity  10Scope: privatex01  INT

11Description: Granularity of size of refined regions

11  Range  10  x01  Default: 1

111:*  10

11minxdistance  10Scope: privatex01  INT

11Description: Minimum distance (in grid points) between coarse and fine grid boundaries

11  Range  10  x01  Default: 4

110:*  10

11

11minxfraction  10Scope: privatex01  REAL

11Description: Minimum fraction of required refined points that need to be present in a refined region

11  Range  10  x01  Default: 0.9

110:*  10

11

11movementxthresholdx1 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx10 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx2 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx3 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx4 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx5 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx6 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx7 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11movementxthresholdx8 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11movementxthresholdx9 10Scope: privatex01  REAL

11Description: Minimum movement to trigger a regridding

11  Range  10  x01  Default: 0.0

110:*  10

11

11numxcentres  10Scope: privatex01  INT

11Description: Number of refinement centres

11  Range  10  x01  Default: (none)

110:10  10

11

11numxlevelsx1  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx10  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx2  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx3  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx4  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx5  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx6  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx7  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11numxlevelsx8  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11numxlevelsx9  10Scope: privatex01  INT

11Description: Number of refinement levels for this centre

11  Range  10  x01  Default: 1

111:30  10

11

11positionxxx1  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx10  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx2  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx3  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx4  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx5  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx6  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx7  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx8  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxxx9  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11positionxyx1  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx10  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx2  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx3  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx4  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx5  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx6  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx7  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx8  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxyx9  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx1  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11positionxzx10  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx2  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx3  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx4  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx5  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx6  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx7  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx8  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11positionxzx9  10Scope: privatex01  REAL

11Description: Position of this centre

11  Range  10  x01  Default: 0.0

11*:*  10

11

11radiusx1  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx10  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11radiusx2  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx3  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx4  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx5  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx6  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx7  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx8  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusx9  10Scope: privatex01  REAL

11Description: Radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10illegal

110:*  10

11

11radiusxrelxchangexthresholdx1 10Scope: privatex01  REAL

11Description: Minimum RELATIVE change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx10 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11radiusxrelxchangexthresholdx2 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx3 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx4 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx5 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx6 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx7 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx8 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxrelxchangexthresholdx9 10Scope: privatex01  REAL

11Description: Minimum change in radius to trigger a regridding

11 Range 10 x01 Default: 0.0
110.0:*  10

11

11radiusxxx1  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx10  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11radiusxxx2  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx3  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx4  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx5  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx6  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx7  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx8  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxxx9  10Scope: privatex01  REAL

11Description: x-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx1  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11radiusxyx10  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx2  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx3  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx4  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx5  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx6  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx7  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx8  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxyx9  10Scope: privatex01  REAL

11Description: y-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11radiusxzx1  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx10  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx2  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx3  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx4  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx5  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx6  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx7  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11radiusxzx8  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11radiusxzx9  10Scope: privatex01  REAL

11Description: z-radius of refined region for this level of this centre

11 Range 10 x01 Default: -1.0
11-1  10ignore this radius

110:*  10

11

11regridxevery  10Scope: privatex01  INT

11Description: Regrid every n time steps

11 Range 10 x01 Default: (none)
11-1  10regrid never
11                       10regrid during initial data calculation only
111:*  10regrid every n time steps

11

11snapxtoxcoarse  10Scope: privatex01  BOOLEAN

11Description: Ensure that the fine grid extent coincides with coarse grid points

11                       10  x01  Default: no

11

11symmetryxparity  10Scope: privatex01  BOOLEAN

11Description: parity

11                       10  x01  Default: no

11

11symmetryxperiodicxx  10Scope: privatex01  BOOLEAN

11Description: Ensure a periodicity symmetry in the x direction

11                       10  x01  Default: no

11

11symmetryxperiodicxy  10Scope: privatex01  BOOLEAN

11Description: Ensure a periodicity symmetry in the y direction

11                       10  x01  Default: no

11

11symmetryxperiodicxz  10Scope: privatex01  BOOLEAN

11Description: Ensure a periodicity symmetry in the z direction

11                       10  x01  Default: no

11

11symmetryxrotating180 10Scope: privatex01  BOOLEAN

11Description: Ensure a 180 degree rotating symmetry about the z axis

11                       10  x01  Default: no

11

11symmetryxrotating90  10Scope: privatex01  BOOLEAN

11Description: Ensure a 90 degree rotating symmetry about the z axis

11                       10  x01  Default: no

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: Display regridding information on the terminal

11                       10  x01  Default: no

11

11veryverbose  10Scope: privatex01  BOOLEAN

11Description: Display much regridding information on the terminal

11                       10  x01  Default: no

7 Interfaces

General Implements: carpetregrid2

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 lastxiteration 10 lastxiteration x01  compact110
11                 10  x01 description11Last iteration at which there was a regridding
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11INT
11 lastxmap 10 lastxmap x01 compact110
11                 10  x01 description11Last map at which there was a regridding at lastxiteration
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11INT
11 oldxactive 10 oldxactive x01 compact110
11                 10  x01 description11Old whether this centre is active
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11INT
11 oldxnumxlevels 10 oldxnumxlevels x01  compact110
11                 10  x01 description11Old number of refinement levels
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11INT
11 oldxpositions 10 x01 compact110
11                 10 oldxpositionxx x01 description11Old positions of refined regions
11                 10 oldxpositionxy x01 dimensions110
11                 10 oldxpositionxz x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL
11 oldxradiixyz 10  x01  compact110
11                 10  oldxradiusxx  x01 description11Old radii of refined regions for each level
11                 10  oldxradiusxy  x01 dimensions111
11                 10  oldxradiusxz  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1130

11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL

11

7.0.2 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 active 10 active x01 compact110
11                 10  x01 description11Whether this centre is active
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11INT
11 numxlevels 10 numxlevels x01 compact110
11                 10  x01 description11Number of refinement levels
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11INT
11  positions  10  x01  compact110
11                 10  positionxx  x01 description11Positions of refined regions
11                 10  positionxy  x01 dimensions110
11                 10  positionxz  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL
11  radii  10  x01  compact110
11                 10  radius  x01 description11Radii of refined regions for each level
11                 10  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1130

11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL
11  radiixyz  10  x01  compact110
11                 10  radiusxx  x01 description11Radii of refined regions for each level
11                 10  radiusxy  x01 dimensions111
11                 10  radiusxz  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1130

11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL
11 levelxmask 10 levelxmask x01 compact110
11                 10  x01 description11Requested refinement level for this grid point
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11 10 x01 tags11checkpoint=”no” prolongation=”none”
11                 10  x01 timelevels111
11                 10  x01variable type11REAL

11

Uses header: Timer.hh bbox.hh bboxset.hh defs.hh dh.hh gh.hh mpixstring.hh region.hh vect.hh carpet.hh

Provides: CarpetxRegrid to CarpetxRegridMaps to 8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetRegrid2. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11Conditional:
11lastxiteration lastxmap  11levelxmask
11active numxlevels positions radii radiixyz  11x
11oldxactive oldxpositions oldxnumxlevels oldxradiixyz11x

11x  11x

11

Scheduled Functions

CCTKxPARAMCHECK x carpetregrid2xparamcheck x check parameters x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxWRAGH x carpetregrid2xinitialise x initialise locations of refined regions x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function
11 x 11Writes: 11carpetregrid2::radius
11 x 11x  11radiixyz
11 x 11x  11carpetregrid2::lastxiteration
11 x 11x  11lastxmap
11 x 11x  11active
11 x 11x  11numxlevels
11 x 11x  11carpetregrid2::positions
11 x 11x  11oldxpositions